What is the ordered pair(x,y) form
3x + 5y = 21
-9x + 4y = -6

Answer by mananth(16949)

3 x + 5 y = 21 .............1
-9 x + 4 y = -6 .............2
Eliminate y
multiply (1)by -4
Multiply (2) by 5
-12 x -20 y = -84
-45 x 20 y = -30
Add the two equations
-57 x = -114
/ -57
x = 2
plug value of x in (1)
3 x + 5 y = 21
6 + 5 y = 21
5 y = 21 -6
5 y = 15
y = 3
(2,3)